OpenStudy (anonymous):

Solve the following inequality:

OpenStudy (anonymous):

\[4 - 2x \le 5 - x + 1\]

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

We will need to move x to one side.

OpenStudy (anonymous):

So do we divide by 2? Or add 2x?

TheSmartOne (thesmartone):

First lets simplify each side.

TheSmartOne (thesmartone):

\[4 - 2x \le 5 - x + 1\] What is \(\sf 5-x+1=?\)

TheSmartOne (thesmartone):

Lets simplify that side so we won't have to do 1 extra step :)

OpenStudy (anonymous):

Ok but I think my teacher wants me to do the extra step lol.

TheSmartOne (thesmartone):

Well the extra step is not necessary. Actualy we are doing this step now so later we won't have have to move 2 numbers over to isolate x.

OpenStudy (anonymous):

Okie

TheSmartOne (thesmartone):

So did you get any value for \(\sf 5-x+1=?\)

OpenStudy (anonymous):

No?

TheSmartOne (thesmartone):

Just simplify it. What is \(\sf 5-x+1=5+1-x=?\)

TheSmartOne (thesmartone):

@purpleowl

OpenStudy (anonymous):

Sorry. I feel really stupid.

TheSmartOne (thesmartone):

What is \(\sf 5+1=?\)

OpenStudy (anonymous):

6

OpenStudy (anonymous):

6 -x ?

TheSmartOne (thesmartone):

So... \(\sf 5-x+1=6-x\)

TheSmartOne (thesmartone):

Correct! :)

OpenStudy (anonymous):

Yay (:

TheSmartOne (thesmartone):

So now we have \(\sf\Large 4 - 2x \le 6-x\)

TheSmartOne (thesmartone):

So lets add 2x to both sides. What do we get?

OpenStudy (anonymous):

\[4 \le 6 + 2x?\]

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

@Legends

OpenStudy (anonymous):

I have to go soon and I have a lot more problems :/ are you guys there @Legends @TheSmartOne

OpenStudy (legends):

Yea sorry i was doing the problem

OpenStudy (anonymous):

Oh okay

TheSmartOne (thesmartone):

Back.

TheSmartOne (thesmartone):

Your answer is wrong. Lets add 2x on both sides. \(\sf\Large 4 - 2x +2x\le 6-x+2x\)

OpenStudy (legends):

@purpleowl Plug in -2 for x and see what you get. 4 - 2x = 6 - x